# Golden Noise for Diffusion Models: A Learning Framework

### v4

@@ Appendix B Related Work
Lugmayr et al. [22] utilizes a pre-trained unconditional diffusion model as a generative prior and alters the reverse diffusion iterations based on the unmasked regions of the input image. Meng et al. [23] observe that denoising the noise with inversion steps can generate better images compared with the original denoising process. Based on that, Qi et al. [27] aims to reduce the truncate errors during the denoising process, by increase the cosine similarity between the initial noise and the inversed noise in an end-to-end way. It introduces significant time costs, and the synthesized images may be over-rendered, making it difficult to use in practical scenarios.
@@ Appendix C Discussion with the Previous Works
We previously mentioned that these methods [3, 7] optimize the noise during the reverse process by incorporating additional modules. These methods have shown promising results in tasks involving compositional generalization. However, these methods often struggle to transfer to other datasets and models, making them not universally applicable. These approaches require the manual interest subject tokens, necessitating extensive test to identify the optimal tokens for a given sentence, which complicates their application across different datasets. Furthermore, modifying the model pipeline usually requires in-depth code changes, making it difficult to achieve straightforward plug-and-play integration with other models. Moreover, these methods demand multiple rounds of noise optimization during the reverse process, resulting in significant time consumption.
In contrast, our approach addresses these challenges from multiple perspectives, offering a more flexible and universal solution. It is capable of cross-model and cross-dataset applications, provides plug-and-play functionality, and incurs minimal time overhead. The experimental results are shown in Table. 7. We follow the code in Guo et al. [7], and manually provide the subject tokens following Chefer et al. [3]. We conduct the experiments on StableDiffusion-v1-4 [30] on Pick-a-Pic dataset. Note that The authors did not provide the dataset or the necessary subject tokens for the algorithm, so we are limited to using our own dataset. Moreover, we had to manually select and trim the objects in the prompts. Additionally, the author’s code requires extensive modifications to the pipeline, making it difficult to adapt for use with other diffusion models. We directly apply the NPNet trained for SDXL to StableDiffusion-v1-4. The results demonstrate the superiority of our NPNet.
@@ D.1 Motivation
To the best of our knowledge, we are the first to propose the noise prompt learning framework, as illustrated in Fig. 6. In order to transform the random Gaussian noise into the golden noise, we utilize the re-denoise sampling, a straightforward method to boost the semantic faithfulness in the synthesized images, illustrated in Fig. 7, to obtain the noise pairs to construct our noise prompt dataset. When designing the architecture of noise prompt network, we discover the singular vectors between the source noise $\mathbf{x}_{T}$ and target noise $\mathbf{x}^{\prime}_{T}$ exhibit remarkable similarity, albeit possibly in opposite directions, as shown in Fig. 8. Building upon this, we design the singular value prediction branch.
@@ D.2 Exploration of Data Selection Strategies
Since the target noise collected through re-denoise sampling is not always of high quality, it is crucial to choose an appropriate method for data filtering. Effective selection ensures that only high-quality noise pairs are used, which is essential for training the NPNet, affecting the model’s performance and reliability. For this reason, we conduct experiments on the choice of human preference model to filter our data, shown in Appendix Table  8, here the filtering threshold $m=0$. The results demonstrate that using HPSv2 ensures data diversity, allowing the filtered data to enhance the model’s performance effectively. This approach helps maintain a rich variety of training samples, which contributes to the model’s generalization ability and overall effectiveness.
@@ D.3 Evaluate the Quality of Synthesized Images
D.4 Generalization and Robustness
@@
D.5 Efficiency Analysis and Ablation Studies
@@
Table 12: The values of the two trainable parameters $\alpha$ and $\beta$.
Model $\alpha$ $\beta$
SDXL 1.00E-04 -0.0189
DreamShaper-xl-v2-turbo 7.00E-05 0.0432
Hunyuan-DiT 2.00E-04 0.0018
@@
D.6 Experiments on Large-scale Dataset
To evaluate the effectiveness of our NPNet, we conduct the experiments on a large-scale dataset, GenEval, across different T2I models. The results are shown in Appendix Table 13. The all improved metrics reveal the superiority of our NPNet, suggesting that our NPNet can improve the compositional image properties of the synthesized images.
@@ Appendix F Discussion
1:  Input: Noise prompt dataset $\mathcal{D}:=\{\mathbf{x}_{T_{i}},\mathbf{x}^{\prime}_{T_{i}},\mathbf{c}_{i}\}_{i=1}^{\lvert\mathcal{D}\rvert}$, noise prompt model $\phi$ parameterized by singular value predictor $f$($\cdot$) and residual predictor $g$($\cdot$, $\cdot$), the frozen pre-trained text encoder $\mathcal{E}$($\cdot$) from diffusion model, normalization layer $\sigma$($\cdot$,$\cdot$ , ⋅), $\mathrm{MSE}$ loss function $\ell$, and two trainable parameters $\alpha$ and $\beta$. 2:  Output: The optimal noise prompt model $\phi^{*}$ trained on the training set $\mathcal{D}$. 13:  # singular value prediction, see equation 6 $\mathbf{\tilde{x}}^{\prime}_{T_{i}}$ = $f$($\mathbf{x}_{T_{i}}$) 24:  # residual prediction, see equation 7 $\mathbf{\hat{x}}_{T_{i}}$ = $g$($\mathbf{x}_{T_{i}}$, $\mathbf{c}_{i}$) 35:  # see equation 9 $\mathbf{x}^{\prime}_{T_{pred_{i}}}$ = $\alpha\sigma(\mathbf{x}_{T_{i}},\mathcal{E}(\mathbf{c}_{i}))$ + $\mathbf{\tilde{x}}^{\prime}_{T_{i}}$ + $\beta\mathbf{\hat{x}}_{T_{i}}$ 6:  $\mathcal{L}_{i}$ = $\ell(\mathbf{x}^{\prime}_{T_{pred_{i}}},\mathbf{x}^{\prime}_{T_{i}})$ 7:  update $\phi$ 8:  return  $\phi^{*}$
@@
Figure 18: Visualization results about different baseline methods on SDXL.

### v5

@@ Appendix B Related Work
Lugmayr et al. [32] utilizes a pre-trained unconditional diffusion model as a generative prior and alters the reverse diffusion iterations based on the unmasked regions of the input image. Hu et al. [19] proposes to denoise one more step before the standard denoising process to eliminate the SNR noise discrepancy between training and inference. Meng et al. [34] observe that denoising the noise with inversion steps can generate better images compared with the original denoising process. Based on that, Qi et al. [38] aims to reduce the truncate errors during the denoising process, by increase the cosine similarity between the initial noise and the inversed noise in an end-to-end way. It introduces significant time costs, and the synthesized images may be over-rendered, making it difficult to use in practical scenarios.
@@ Appendix C Discussion with the Previous Works
We previously mentioned that these methods [5, 11, 16, 17, 47], optimize the noise during the reverse process by incorporating additional modules. These methods have shown promising results in tasks involving compositional generalization. However, these methods often struggle to transfer to other datasets and models, making them not universally applicable. These approaches require the manual interest subject tokens, necessitating extensive test to identify the optimal tokens for a given sentence, which complicates their application across different datasets. Furthermore, modifying the model pipeline usually requires in-depth code changes, making it difficult to achieve straightforward plug-and-play integration with other models. Moreover, these methods demand multiple rounds of noise optimization during the reverse process, resulting in significant time consumption.
In contrast, our approach addresses these challenges from multiple perspectives, offering a more flexible and universal solution. It is capable of cross-model and cross-dataset applications, provides plug-and-play functionality, and incurs minimal time overhead. We first follow the code in Guo et al. [11], and manually provide the subject tokens following Chefer et al. [5]. We conduct the experiments on StableDiffusion-v1-4 [41] on Pick-a-Pic dataset. Note that The authors did not provide the dataset or the necessary subject tokens for the algorithm, so we are limited to using our own dataset. Moreover, we had to manually select and trim the objects in the prompts. Additionally, the author’s code requires extensive modifications to the pipeline, making it difficult to adapt for use with other diffusion models. We directly apply the NPNet trained for SDXL to StableDiffusion-v1-4. The experimental results are shown in Table. 7, demonstrating the superiority of our NPNet. In addition, to prove that our method can be directly used in conjunction with other noise optimization methods, we directly use the NPNet trained on SDXL on these mainstream noise optimization methods, and the experimental results in Table. 20 and Fig. 19 prove that our NPNet can further improve the performance of other methods, which validates the generalizability of our NPNet.
Table 8: We evaluate the effectiveness of our NPNet on T2I-CompBench benchmark. The results validate the effectiveness of our method.
Attribute Binding Object Relationship
Method Color (↑) Shape (↑) Texture (↑) 2D-Spatial (↑) 3D-Spatial (↑) Non-Spatial (↑) numeracy (↑) Complex (↑)
Standard 0.5850 0.5028 0.5083 0.2097 0.3667 0.3113 0.5013 0.3130
Ours 0.5940 0.5094 0.5207 0.2146 0.3713 0.3151 0.5201 0.3174
@@ D.1 Motivation
To the best of our knowledge, we are the first to propose the noise prompt learning framework, as illustrated in Fig. 6. In order to transform the random Gaussian noise into the golden noise, we utilize the re-denoise sampling, a straightforward method to boost the semantic faithfulness in the synthesized images, illustrated in Fig. 7, to obtain the noise pairs to construct our noise prompt dataset. Notably, in Fig. 9, we validate that re-denoising sampling can enhance the semantic information in the initial noise. When designing the architecture of noise prompt network, we discover the singular vectors between the source noise $\mathbf{x}_{T}$ and target noise $\mathbf{x}^{\prime}_{T}$ exhibit remarkable similarity, albeit possibly in opposite directions, as shown in Fig. 8. Building upon this, we design the singular value prediction branch.
@@ D.2 Exploration of Data Selection Strategies
Since the target noise collected through re-denoise sampling is not always of high quality, it is crucial to choose an appropriate method for data filtering. Effective selection ensures that only high-quality noise pairs are used, which is essential for training the NPNet, affecting the model’s performance and reliability. For this reason, we conduct experiments on the choice of human preference model to filter our data, shown in Appendix Table  9, here the filtering threshold $m=0$. The results demonstrate that using HPSv2 ensures data diversity, allowing the filtered data to enhance the model’s performance effectively. This approach helps maintain a rich variety of training samples, which contributes to the model’s generalization ability and overall effectiveness.
@@ D.3 Evaluate the Quality of Synthesized Images
D.4 Evaluate the probability density of synthesized images with NPNet
We visualize image distributions before and af- ter NPNet by generating 50 images under identical conditions, using ImageNet images (1 class) as ground truth and text-to- image SDXL for generation in Fig. 13. The experimental results show the use of NPNet, which enables the generated images to more closely resemble the real distribution, and more inclined to high-density reigons.
Figure 13: We visualize image distributions before and after NPNet by generating 50 images under identical conditions, using ImageNet images (1 class) as ground truth and text-to-image SDXL for generation. While text-to-image (v.s. class-conditioned) generation may induce distribution shifts, NPNet shifts generation toward high-density regions.
D.5 Generalization and Robustness
@@
D.6 Efficiency Analysis and Ablation Studies
@@
D.7 Experiments on Large-scale Dataset
To evaluate the effectiveness of our NPNet, we conduct the experiments on a large-scale dataset, GenEval, across different T2I models. The results are shown in Appendix Table 13 and Table 14. The all improved metrics reveal the superiority of our NPNet, suggesting that our NPNet can improve the compositional image properties of the synthesized images.
@@ D.7 Experiments on Large-scale Dataset
Table 14: We evaluate the effectiveness of our NPNet on Genval benchmark.
Single (↑) Two (↑) Counting (↑) Colors (↑) Positions (↑) Color Attribution (↑) Overall (↑)
Standard 97.50 63.64 36.25 85.11 8.00 19.00 51.58
NPNet (ours) 98.75 67.68 37.50 82.98 13.00 21.00 53.49
@@ Appendix F Discussion
Table 15: The values of the two trainable parameters $\alpha$ and $\beta$.
Model $\alpha$ $\beta$
SDXL 1.00E-04 -0.0189
DreamShaper-xl-v2-turbo 7.00E-05 0.0432
Hunyuan-DiT 2.00E-04 0.0018
@@
Table 18: We evaluate NPNet on few steps T2I diffusion models, like LCM [33], PCM [51] and SDXL-Lightning [26] on GenEval dataset. Here we use the NPNet from SDXL, and the results demonstrate NPNet can generalize well to different kinds of T2I diffusion models, boosting their performance directly.
Model PickScore (↑) HPSv2 (↑) AES (↑) ImageReward (↑) CLIPScore (↑)
Standard 22.85 29.12 5.6521 59.02 0.8093
SDXL-Lightning (4 steps) NPNet(ours) 23.03 29.71 5.7178 72.67 0.8150
Standard 22.30 26.52 5.4932 33.21 0.8050
LCM (4 steps) NPNet(ours) 22.38 26.83 5.5598 37.08 0.8123
Standard 22.05 26.98 5.5245 23.28 0.8031
PCM (8 steps) NPNet(ours) 22.22 27.59 5.5667 35.01 0.8175
@@
1:  Input: Noise prompt dataset $\mathcal{D}:=\{\mathbf{x}_{T_{i}},\mathbf{x}^{\prime}_{T_{i}},\mathbf{c}_{i}\}_{i=1}^{\lvert\mathcal{D}\rvert}$, noise prompt model $\phi$ parameterized by singular value predictor $f$($\cdot$) and residual predictor $g$($\cdot$, $\cdot$), the frozen pre-trained text encoder $\mathcal{E}$($\cdot$) from diffusion model, normalization layer $\sigma$($\cdot$,$\cdot$ , ⋅), $\mathrm{MSE}$ loss function $\ell$, and two trainable parameters $\alpha$ and $\beta$. 2:  Output: The optimal noise prompt model $\phi^{*}$ trained on the training set $\mathcal{D}$. 13:  # singular value prediction, see (6) $\mathbf{\tilde{x}}^{\prime}_{T_{i}}$ = $f$($\mathbf{x}_{T_{i}}$) 24:  # residual prediction, see (7) $\mathbf{\hat{x}}_{T_{i}}$ = $g$($\mathbf{x}_{T_{i}}$, $\mathbf{c}_{i}$) 35:  # see (9) $\mathbf{x}^{\prime}_{T_{pred_{i}}}$ = $\alpha\sigma(\mathbf{x}_{T_{i}},\mathcal{E}(\mathbf{c}_{i}))$ + $\mathbf{\tilde{x}}^{\prime}_{T_{i}}$ + $\beta\mathbf{\hat{x}}_{T_{i}}$ 6:  $\mathcal{L}_{i}$ = $\ell(\mathbf{x}^{\prime}_{T_{pred_{i}}},\mathbf{x}^{\prime}_{T_{i}})$ 7:  update $\phi$ 8:  return  $\phi^{*}$
@@
Figure 19: Winning rate of noise optimization methods on three different datasets, using SDXL. The results demonstrate that when NPNet is used in conjunction with other noise optimization methods, it can enhance the winning rate of existing approaches to some extent.
Table 20: We combine other popular noise optimization methods with our NPNet, evaluating on three different datasets on SDXL. The experimental results indicate that when NPNet is used alongside other methods, it significantly enhances their performance, further validating the effectiveness and generalizability of our approach.
Dataset Method PickScore (↑) HPSv2 (↑) AES (↑) ImageReward (↑) CLIPscore (↑)
Standard 21.69 28.48 6.0373 58.01 0.8204
Re-sampling [32] 21.77 28.63 5.9875 64.94 0.8327
+ NPNet (ours) 21.90 29.29 6.1491 71.09 0.8386
PAG [1] 21.64 29.45 6.2246 55.91 0.7966
+ NPNet (ours) 21.70 29.80 6.2411 62.03 0.8079
CFG++ [9] 21.67 29.54 6.1239 75.16 0.8322
+ NPNet (ours) 21.82 29.84 6.1703 81.60 0.8374
APG [44] 21.69 28.48 6.1472 65.86 0.8295
+ NPNet (ours) 21.86 29.13 6.1629 76.50 0.8322
FreeU [47] 21.39 29.12 6.2134 79.74 0.8094
+ NPNet (ours) 21.42 29.44 6.2194 77.42 0.8059
SAG [18] 21.70 29.42 6.1507 59.61 0.8162
+ NPNet (ours) 21.79 29.63 6.1535 65.75 0.8193
SEG [16] 21.47 29.23 6.2118 61.60 0.8060
Pick-a-Pic + NPNet (ours) 21.60 29.72 6.2253 72.51 0.8077
Standard 22.31 26.72 5.5952 62.21 0.8077
Re-sampling 22.30 26.96 5.5104 64.07 0.8106
+ NPNet (ours) 22.47 27.55 5.6487 75.36 0.8136
PAG 22.25 27.81 5.7082 67.25 0.7965
+ NPNet (ours) 22.34 28.17 5.7646 72.10 0.7989
CFG++ 22.35 28.28 5.6720 81.69 0.8230
+ NPNet (ours) 22.45 28.67 5.7035 86.40 0.8263
APG 22.29 26.94 5.6180 70.15 0.8194
+ NPNet (ours) 22.46 27.73 5.6445 81.30 0.8182
FreeU 21.97 27.31 5.7379 63.18 0.7973
+ NPNet (ours) 22.04 27.87 5.7464 77.53 0.7998
SAG 22.30 27.64 5.6651 64.56 0.8144
+ NPNet (ours) 22.37 27.98 5.6998 72.18 0.8104
SEG 22.16 28.08 5.7867 64.37 0.7979
DrawBench + NPNet (ours) 22.30 28.51 5.8093 74.12 0.7985
Standard 22.88 29.71 5.9985 96.63 0.8734
Re-sampling 22.91 29.78 5.9948 97.39 0.8775
+ NPNet (ours) 22.96 30.16 6.0098 98.34 0.8787
PAG 22.80 30.54 6.1180 90.94 0.8536
+ NPNet (ours) 22.92 30.97 6.1091 106.24 0.8584
CFG++ 23.03 30.94 6.0269 107.70 0.8875
+ NPNet (ours) 23.08 31.07 6.0533 109.06 0.8920
APG 22.94 29.64 6.0572 97.90 0.8836
+ NPNet (ours) 23.06 30.40 6.0232 111.55 0.8945
FreeU 22.72 30.52 6.0907 97.03 0.8580
+ NPNet (ours) 22.73 30.83 6.1148 102.77 0.8631
SAG 22.86 30.37 6.0680 93.62 0.8670
+ NPNet (ours) 22.92 30.71 6.0177 97.20 0.8748
SEG 22.74 30.47 6.0841 91.86 0.8569
HPD + NPNet (ours) 22.80 30.74 6.1160 104.16 0.8591
Appendix G More Visualization Results
We present more visualization results on different dataset, different diffusion models and with different noise optimization methods on Fig. 22,  23,  24,  25,  26,  27,  28,  29,  31,  30,  32 and  33.
@@ Appendix G More Visualization Results
Figure 23: Visualization results about different methods on SDXL.
Figure 24: Qualitative comparison on LCM.
Figure 25: Qualitative comparison on PCM.
Figure 26: Qualitative comparison on SDXL-Lightning.
Figure 27: Qualitative comparison on Re-sampling.
Figure 28: Qualitative comparison on PAG.
Figure 29: Qualitative comparison on CFG++.
Figure 30: Qualitative comparison on FreeU.
Figure 31: Qualitative comparison on APG.
Figure 32: Qualitative comparison on SAG.
Figure 33: Qualitative comparison on SEG.